OpenStudy (anonymous):

One exterior angle of a regular polygon measures 40. What is the sum of the polygon's interior angle measures?

ganeshie8 (ganeshie8):

Use this :- Sum of exterior angles add up to 360, for a regular n sided polygon : n * angle = 360 n * 40 = 360 n = 9 so its a 9 sided polygon.

ganeshie8 (ganeshie8):

Interior angle = 180 - exterior angle = 180 - 40 = 140

OpenStudy (anonymous):

so its 9 x 140 ??

ganeshie8 (ganeshie8):

thats perfect ! you got it !!

OpenStudy (anonymous):

so my answer would be 1260 ?

ganeshie8 (ganeshie8):

yep

OpenStudy (anonymous):

thank you :D